OpenStudy (anonymous):

Write an equation in slope-intercept form for the line perpendicular to y=1/2x +7 that passes through (-1,3)

OpenStudy (ghazi):

use y-y1= m(x-x1) and here m= negative inverse of the slope of given line...since required line is perpendicular to the given therefore it's (y-3)= -2(x+1)

OpenStudy (anonymous):

wait quick other question sorry lol (14-a^0b^2)^-3 subsitute a=-2 and b=4 I know how to subsitute, but i'm confused about the negatives. when you subsitute the -2, is it this? and does the 2 become -1 or 1?